Macy's father will help her buy a house for $100,000 with a 10 - year loan. He does not mind lending her money at the same rate that his money market account provides, which is 6% per year. Macy feels she can give her dad $25,000 at the end of 10 years from her savings account while maing quareterly payments. What quareterly amount, paid at the end of every quarter, will shde have to give to her dad?

What would the quarterly payments be if made at the beginning of the quarter?
Answer #1

At the end of 10th year, Macy's Father will receive $100,000 via

1. One time payment of $25000

2. Quarterly payments over 10 years that will amount to $75000 at the 10th year

To Calculate Quarterly Amount,

N = 10*4 (Since it is quarterly payment), I = 6/4 = 1.5%(Quarterly Payment), PV = 0, FV = 75000

=PMT(1.5%, 40, 0, 75000)

Installment per quarter = $1382.03